The Web-based company Oh Baby! Gifts has a goal of processing 95 percent of its orders on the same day they are received. If the 485 out of the next 500 orders are processed on the same day, would this prove that they are exceeding their goal, using α = .025?
Answer by stanbon(57307)

The Web-based company Oh Baby! Gifts has a goal of processing 95 percent of its orders on the same day they are received. If the 485 out of the next 500 orders are processed on the same day, would this prove that they are exceeding their goal, using α = .025?
Ho: p <= 0.95
Ha: p > 0.95
-----------------
sample proportion: phat = 485/500 = 0.97
-----
Test Stat: t(0.97) = (0.97-0.95)/sqrt[0.95*0.05/500] = 2.0520
--------
p-value = P(z > 2.0520) = 0.02001
---
Conclusion: Since p-value is less than alpha, reject Ho.
The test gives some (not much) evidence at the 2.5% significance
level that the goal might be being exceeded.
================================================
Cheers,
Stan H.